OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + y = 1 6x + 2y = −9 0 2 There is no x value as there is no solution to this system x can be any value as there are infinitely many solutions to this system

OpenStudy (mathmate):

Subtract twice the first equation from the second. If you get 0=-11, then there is no solution. If you get 0=0, then there are infinitely many solutions.

OpenStudy (anonymous):

These equations form two lines. -If the slope is different then they will at some point meet each other. -if the slope is the same they will be parrallel and will: a) never touch b) always touch

OpenStudy (mathstudent55):

Multiply the first equation by 2: 3x + y = 1 2(3x + y) = 2(1) 6x + 2y = 2 The second equation is 6x + 2y = -9 If 6x + 2y = 1 and 6x + 2y = 9, then the conclusion is 1 = -9, which is a false statement. Therefore there is no solution.

OpenStudy (anonymous):

ax + by +c = 0 slope = -a/b

OpenStudy (anonymous):

Never touch??